OpenStudy (clamin):

HELP!! MEDAL!! What value is missing from the table?? a) 8.5 b) 6 c) 5 d) 7.5

OpenStudy (anonymous):

what table

OpenStudy (clamin):



OpenStudy (anonymous):

i think 6

OpenStudy (clamin):

how do you ssolve it?? @limecake08

OpenStudy (kaylardz0405):

have you found any patterns

OpenStudy (clamin):

no

OpenStudy (anonymous):

what do u have a clue of what it will be itwas dumm of me giving direct

OpenStudy (kaylardz0405):

im really not sure but i noticed that the more speed the less time the less speed the more time

OpenStudy (kaylardz0405):

so think about that

OpenStudy (kaylardz0405):

i think maybe 5

OpenStudy (anonymous):

look frome the bottom to the top and dont give direct

OpenStudy (kaylardz0405):

yeah im pretty sure of 5

OpenStudy (anonymous):

u are right cus it says 3.75

OpenStudy (anonymous):

so can i have a medal and the answer is 5 i just explaned it

OpenStudy (triciaal):

the speed times the time is constant (distance)

OpenStudy (anonymous):

so it is d or c ...........................c

OpenStudy (kaylardz0405):

but its defintly not A or B because the whole nubers have a 5 at the end and the . have other nubers

OpenStudy (clamin):

now im confuse

OpenStudy (triciaal):

vt = d d = 75 v = 15 find t t = d/v = 75/15 =

OpenStudy (anonymous):

ok the answer is cccccccccccccccccccc

OpenStudy (anonymous):

sorry me ans @KaylaRdz0405 just took alot of time explaning it

OpenStudy (anonymous):

and

OpenStudy (triciaal):

@clamin do you understand the process now? still have questions?

OpenStudy (anonymous):

medal please

OpenStudy (clamin):

now i get it thanks

OpenStudy (triciaal):

welcome